OpenStudy (anonymous):

A sample of chromium weighing 465 g was initially at a temperature of 26.45 degrees Celsius. It required 1027 joules of heat energy to increase the temperature to 31.35 degrees Celsius. 1. What is the specific heat capacity of the Chromium 2. What is the molar heat capacity of the Chromium 3. What is the heat capacity of the Chromium

OpenStudy (joannablackwelder):

Specific heat capacity = q/(mΔT) Molar heat capacity = specific heat capacity * molar mass Heat capacity = q/ΔT where q is energy, m is mass, and ΔT is change in temp (in deg C or K)

OpenStudy (anonymous):

Thank You